The length of a rectangle is 4 cm longer than the with. If the perimeter is 16 cm, find the width And length of the rectangle
Answer by algebrahouse.com(1659)

Perimeter of a rectangle
= 2(width) + 2(length)

w = width
w + 4 = length {length is 4 longer than width}

2w + 2(w + 4) = 16 {substituted into perimeter formula}
2w + 2w + 8 = 16 {used distributive property}
4w + 8 = 16 {combined like terms}
4w = 8 {subtracted 8 from each side}
w = 2 {divided each side by 4}
w + 4 = 6 {substituted 2, in for w, into w + 4}

width = 2 cm
length = 6 cm

For more help from me, visit: www.algebrahouse.com